# references.bib

### v1

@@ 1 Introduction
The need for diagnostic systems to help pathologists manage the anticipated workload increases as cancer cancers worldwide are on the rise [moscalu_histopathological_2023]. As [sung_global_2021] estimate, breast, colorectal, prostate, and lung cancers are among the six most common cancers types. Projections suggest that cases of these cancers will continue to increase, posing significant challenges due to time-consuming diagnosis, increased demand for tumor subtyping, and personalized treatment [rahib_estimated_2021, soerjomataram_planning_2021, siegel_cancer_2023]. Deep learning (DL) has made significant progress in medical imaging, particularly in the field of computational pathology (CPath). Some studies have demonstrated that DL models even surpass human performance in certain tasks, making DL models effective tools to help pathologists cope with the increasing workload [zhang_pathologist-level_2019, liu_deep_2020]. However, the unavailability of the required large data sets and the needed investment of time and effort limits the effectiveness and impact of DL models in pathology.
Recent advances in self-supervised learning have enabled the training of deep neural networks on large amounts of unlabeled medical data, resulting in the creation of foundation models in computer vision [zhang_challenges_2023]. These models are pre-trained on a wide range of images, primarily using self-supervision through contrastive learning or masked image modeling. They have been shown to perform well in downstream tasks, including patch classification and weakly labeled whole slide image (WSI) classification [wang_transformer-based_2022, campanella_computational_2023]. Projects such as the Tissue Cancer Genome Atlas Program (TCGA) provide a data source of thousands of WSIs for training these networks on real-world data. This vast amount of data is necessary for self-supervised trained networks to reach their full potential [zhang_challenges_2023]. However, the amount of resources required to create, train, and deploy such models has raised concerns among researchers about the environmental and other impacts [dhar_carbon_2020, kaack_aligning_2022, the_lancet_digital_health_curbing_2023]. In addition, extended training periods of several weeks impede development cycles and prolong research time.
@@
This paper demonstrates that training a foundation model on supervised signals in CPath using MTL requires less data, time, and energy compared to models trained with self-supervision. At the same time, the measured performance is similar to that obtained from models trained on about 17 times more data without supervision. Following the MTL training scheme presented in Figure 1a, this paper presents Tissue Concepts (TC), a robust encoder that is trained on a mixture of diverse annotations from small and medium-sized datasets in CPath to learn different concepts related to tissue. Considering the future prediction of cancer cases and clinical workflow, we evaluated the performance of the encoder on the four major cancer types breast, colon, lung, and prostate, for whole slide image classification, as shown in Figure 1b. In addition, since models trained on one site are known to perform worse when evaluated on different sites, we test the performance of Tissue Concepts using a cross-center evaluation scheme [howard_impact_2021].
The paper’s main contributions can be summarized as follows.
@@ 2 Related Work
First approaches using MTL in CPath were presented by[mormont_multi-task_2021] and [graham_one_2023]. Mormont and colleagues converted different datasets into 22 classification tasks to train a shared network and contrasted the learned encoders against ImageNet-weights. A SVM was trained on the latent representations of the encoder. They found that the representations perform equal or better to the baseline ImageNet-weights. Graham et al. then used MTL on segmentation and classification tasks. This research focused on specific tasks that were present in the pre-training. However, the evaluation of the general-purpose encoder and the corresponding latent representations based on whole slide image classification combined with cross-center evaluation is still an unexplored area. In addition, general purpose encoders in the form of foundation models have not been considered by [graham_one_2023].
In [schafer_overcoming_2023] we presented a first approach using MTL to train supervised foundation models. By utilized expert knowledge in the form of multi-task learning we trained a shared model, called UMedPT, which can be applied to various medical images. To achieve this, different imaging domains, such as CT, X-ray, and microscopic images were used to train a shared backbone on classification, segmentation, and detection tasks. Currently, the impact of tasks outside the histopathology domain remains unclear due to the diverse pre-training of the encoder. This impact on performance and robustness requires further investigation.
The following sections focus in more detail on two topics discussed in this paper. While foundation models are still largely unexplored in terms of their application and performance, some approaches are mentioned below.
@@ 2.1 Foundation Models
[wang_transformer-based_2022] used data from the TCGA in combination with data from the pathology AI platform PAIP to train a modified swin transformer, called CTransPath (CTP), using self-supervision on 15 million patches. They presented an adapted contrastive loss, based on MoCo v3 [chen_empirical_2021], which uses a memory bank to retrieve the top S semantically relevant entries. These entries were used as additional positive examples for the loss calculation. The authors evaluated their model using patch classification, image retrieval, and weakly labeled WSI classification. Due to the large number of training images and the slow convergence of self-supervised training, they reported a training time of 250 hours on 48 GPUs (12.000 GPU-hours).
[campanella_computational_2023], presented a comparable approach, training a tiny vision transformer (ViT) using standard DINO, and a ViT base model using standard masked autoencoding (MAE), both trained on about 3 billion patches. The authors report a training time of over 3000 GPU-hours for the models that were evaluated on a variety of tasks ranging from disease detection to outcome prediction. The evaluation also included images scanned at a different hospital than the training slides.
[chen2024towards] present a general purpose foundation model that leverages over 100 million patches from 100.000 WSI slides across 20 major cancer types. They train a large ViT on patches, that were collected in an internal dataset, using DINOv2 [oquab2023dinov2]. They evaluate the model on 34 Tasks and find that it surpasses the previous baselines on most of them. The model was trained on 24 80GB GPUs.
Overall, all of the presented models rely on large image databases and require long training times, which contributes to increased CO2 emissions. The presented TC encoder and MTL training aim to reduce the need for large amounts of data while maintaining the desired performance. In addition cross-center evaluation is needed to accurately asses models’ performances.
@@ 2.2 Weakly Labeled WSI Classification
Learning from WSIs that are only labeled on a case basis, or that have only one endpoint, is challenging because training on the entire image at once typically exceeds the GPU memory. In addition, since a WSI provides only one sample, many WSIs are needed to effectively train a deep learning model. Classification of such gigapixel images is therefore typically performed using multiple instance learning (MIL) [campanella_clinical-grade_2019, ghaffari_laleh_benchmarking_2022]. Using MIL involves two parts: first, extracting features from patches of the WSI using a pre-trained encoder to convert them into their latent representations, and second, aggregating features from a WSI using a trainable MIL head to predict the given label [liu_advmil_2024]. Therefore, robust encoders are needed to obtain patch representations that facilitate the second step of MIL [godson_immune_2024]. In this paper MIL is used as an evaluation procedure to test the representativeness of the encoder’s features. The following presents the most commonly used approaches, that focus on solving the second stage of MIL using either attention or convolution-based methods.
[lu_data-efficient_2021] introduced CLAM, a clustering-constrained attention MIL algorithm. The authors trained an attention-based head on features extracted from patches of a WSI to classify the corresponding labels. The attention was then used to identify sub-regions of high diagnostic value, which in turn were used to classify the entire slide. In addition, instance-level clustering was applied over the representative regions to constrain and refine the feature space.
@@ 3.2 Encoder Training
Since transformer-based approaches are currently dominant in the literature we chose a tiny swin transformer as shared encoder [liu_swin_2021, azad_advances_2024] (further denoted as TC-swin). Additionally, we compared the attention-based backbone to a tiny ConvNeXt architecture that shares a similar number of parameters but follows a convolution-based approach [liu_convnet_2022] (further denoted as TC-conv). Both networks were initialized with corresponding ImageNet-1k weights which were imported from the torchvision library [russakovsky_imagenet_2015]. During pre-training, the networks were optimized for 5 million steps until convergence, where one step was defined as computing the loss for one batch of one task. The AdamW optimizer [loshchilov_decoupled_2019] was used with a constant learning rate of $10^{-4}$, a weight decay of 0.01, and gradient accumulation of 128 steps. In total, the models were trained for 160 hours on one NVidia RTX A5000 GPU and the final models were used for evaluation.
@@ 4.2 Evaluation datasets
The learned representations of Tissue Concepts were evaluated using weakly labeled WSI classification tasks, colorectum, represent tissues within the training domain, lung cancer tissue was deliberately chosen to represent tissues outside the training domain. The architecture and learning schedule as described in Section 4.3, was used during the training of each MIL head. Where possible, multi-site datasets were split so that evaluation sites were not included in the training. The feature extractors were evaluated on the following datasets for each organ:
@@
Prostate Cancer: The Prostate cANcer graDe Assessment (PANDA) challenge provides 10616 openly available WSIs from two different centers, Karolinska and Radboud [bulten_artificial_2022]. Each of the WSIs was assigned an ISUP score resulting from the primary and secondary Gleason pattern. The score ranges from 0 to 5, with 0 representing normal tissue and higher numbers representing more severe prostate cancer. This information was used to create binary classification task for evaluation, where the task consists of differentiating between normal (ISUP 0) and cancerous tissue (ISUP > 0). To test cross-center performance and overall performance, three different experiments were created from this dataset. First, the entire set of images obtained from the Radboud UMC was used as both the training and validation set, using a split ratio of 90% for training and 10% for validation. All images provided by the Karolinska Institute were used as the hold-out test set. In a subsequent experiment, the roles of training and test centers were swapped. In addition, a 5-fold cross-validation experiment involving both centers was conducted. We note that a small percentage of the dataset in the form of pre-selected patches were used during pre-training of the encoder, however, the task-specific head was discarded. [mahdi_behzadi_weakly-supervised_2022] used this dataset in a k-fold cross-validation to evaluate the performance of their model and reported the accuracy and area under the curve for binary classification. Their reported results serve as an external baseline. [mahdi_behzadi_weakly-supervised_2022] reported the accuracy and AUC of their algorithm on the dataset. To be comparable, we also report these metrics in our experiments. Additionally, we report the F1-score as per-class metric, as suggested by [maier-hein_metrics_2024].
@@
Colorectal Cancer: During the SemiCOL challenge, a dataset of 499 WSIs from 4 different centers was provided as training dataset. The challenge task was to predict whether a given slide contained cancerous tissue or normal tissue. During the challenge, we used a modified model presented in [schafer_overcoming_2023] as an initial starting point and further fine-tuned on several colon and colorectal tasks. Using the learned encoder to extract features from the WSI, a MIL head was successively trained on these features. This resulted in an external test AUC of 0.99. In the experiments presented in this paper, the challenge training data was split by slide provider, effectively creating four smaller subsets of the original challenge dataset. For the evaluation presented in this paper, four different experiments were conducted. Images from one center were used as training examples (90/10), while images from the other three centers served as the hold out test set. By treating each center as training provider once, and using the other three centers as test centers, four different configurations were created We note that a small percentage of patch data of the dataset was used during pre-training of the encoder, however, the task-specific head was discarded. To the best of our knowledge, there is no other publication reporting results on this specific dataset, so we are unable to report an external baseline. Since the dataset is balanced with regard to the class distribution for each center in all splits, we reported the Accuracy, F1-Score and AUC as metrics, following [maier-hein_metrics_2024].
@@ 4.3 Training of the MIL head
To aggregate the latent WSIs, a small CNN with global max pooling, was used. The initial layer utilized a 1x1 convolution, reducing the latent WSIs’ 768 channel to 16. Following this, 3x3 convolutions with a stride of 2, and padding of 1 doubling in depth each time were applied. Instance normalization, leaky ReLU activation, and 10% 2D dropout were applied in the convolutional layers. The lowest feature pyramid of 64 channels was globally max pooled, followed by one hidden linear layer and a classification layer. This resulted in 19250 total trainable parameters. The convolutional layers were initialized by drawing from a Kaiming normal distribution with a fixed seed. Additionally, we tested an attention-based approach by replacing the max pooling layer with an attention pooling layer, as suggested by [ilse2018attention]. Overall, we found similar results with both approaches across all organs. The results using an ABMIL aggregation head can be found in the subsection B.1.)
@@ 5.2 Prostate Cancer
Table 3: Prostate: Tissue-specific model performance when applied to the PANDA challenge data. The model consist of organ-specific head and frozen encoders: ImageNet, CTP, TC-Swin, TC-Conv. The table shows mean weighted Accuracy, F1 score, AUC, and standard deviations over 5 different runs for 5-fold cross-validation and cross-center transfer performance. The best score is shown in bold.
@@
Table 3 shows the results obtained of five runs on the PANDA dataset with cross-center and mixed evaluation. The ’clinic’ column indicates different settings, where "Karolinska $\rightarrow$ Radboud" refers to the MIL head being trained on data from the Karolinska Institute and evaluated on data from the Radboud Clinic.
When comparing the results obtained in the 5-fold cross-validation with the external baseline, both TC encoders show performances that were close to the external baseline, with TC-conv slightly surpassing it with a mean accuracy of 0.957. The mean accuracies of CTP and ImageNet are 2 and 5 percentage points below the selected baseline, respectively. All foundation models outperform the pure ImageNet encoder by at least three percentage points. In terms of accuracy and AUC scores, both TC encoders slightly outperform the CTP weights by one percentage point with mean acc of 0.936 and mean auc of 0.979.
In the cross-center evaluation, the performance of all models decreases compared to an in-center evaluation. CTP achieves a mean accuracy of 0.819 when transferring decision boundaries from the Karolinska site to Radboud subset. The opposite scenario results in a decrease of 14 percentage points (0.678 mean acc). Furthermore, an even steeper drop in mean accuracy and AUC from 5-fold cross-validation to cross-center validation is observed when using the ImageNet weights (from 0.900 to 0.658). Both TC encoders are able to extract robust features. The TC-Swin encoder (from 0.942 to 0.846) shows more stable performance than the TC-Conv encoder (from 0.957 to 0.764) when compared to the metrics reported in the cross-center evaluation.

### v2

@@ 1 Introduction
The need for diagnostic systems to help pathologists manage the anticipated workload increases as cancer cancers worldwide are on the rise [moscalu_histopathological_2023]. As [sung_global_2021] estimate, breast, colorectal, prostate, and lung cancers are among the six most common cancers types. Projections suggest that cases of these cancers will continue to increase, posing significant challenges due to time-consuming diagnosis, increased demand for tumor subtyping, and personalized treatment [rahib_estimated_2021, soerjomataram_planning_2021, siegel_cancer_2023]. Deep learning (DL) has made significant progress in medical imaging, particularly in the field of computational pathology (CPath). Some studies have demonstrated that DL models even surpass human performance in certain tasks, making them effective tools to help pathologists cope with the increasing workload [zhang_pathologist-level_2019, liu_deep_2020]. However, the unavailability of the required large data sets and the needed investment of time and effort limits the effectiveness and impact of DL models in pathology.
Recent advances in self-supervised learning have enabled the training of deep neural networks on large amounts of unlabeled medical data, resulting in the creation of foundation models in computer vision [zhang_challenges_2023]. These models are pre-trained on a wide range of images, primarily using self-supervision through contrastive learning or masked image modeling. They have been shown to perform well in downstream tasks, including patch classification, and weakly labeled whole slide image (WSI) classification [wang_transformer-based_2022, campanella_computational_2023]. Projects such as the Tissue Cancer Genome Atlas Program (TCGA) provide an openly available data source of thousands of WSIs for training these networks on real-world data. This vast amount of data is necessary for self-supervised trained networks to reach their full potential [zhang_challenges_2023]. However, the amount of resources required to create, train, and deploy such models has raised concerns among researchers about environmental and other impacts [dhar_carbon_2020, kaack_aligning_2022, the_lancet_digital_health_curbing_2023]. In addition, extended training periods of several weeks impede development cycles and prolong research time.
@@
This paper demonstrates that training a foundation model on supervised signals in CPath using MTL requires less data, time, and energy compared to models trained with self-supervision. At the same time, the measured performance is similar to that obtained from models trained on about 17 times more data without supervision. Following the MTL training scheme presented in Figure 1a, this paper presents Tissue Concepts (TC), a robust encoder that is trained on a mixture of diverse annotations from small and medium-sized datasets in CPath to learn different concepts related to tissue. Considering the future prediction of cancer cases and clinical workflow, we evaluated the performance of the encoder on the four major cancer types, breast, colon, lung, and prostate, for classification of entire slide images, as shown in Figure 1b. In addition, since models trained on one site are known to perform worse when evaluated on different sites, we tested the performance of Tissue Concepts using a cross-center evaluation scheme [howard_impact_2021].
The main contributions of the paper can be summarized as follows.
@@ 2 Related Work
First approaches using MTL in CPath were presented by[mormont_multi-task_2021] and [graham_one_2023]. Mormont and colleagues converted different datasets into 22 classification tasks to train a shared network and contrasted the learned encoders against ImageNet weights. The latent representations of the encoder were used to train an SVM. They found that the representations perform equally or better than the baseline ImageNet weights. Graham et al. then used MTL on segmentation and classification tasks. This research focused on specific tasks that were present in the pre-training. However, the evaluation of the general-purpose encoder and the corresponding latent representations based on whole slide image classification combined with cross-center evaluation is still an unexplored area. In addition, general purpose encoders in the form of foundation models have not been considered by [graham_one_2023].
In [schafer_overcoming_2023], we presented a first approach using MTL to train supervised foundation models. Using expert knowledge in the form of multi-task learning we trained a shared model, called UMedPT, which can be applied to various medical images. To achieve this, different imaging domains, such as CT, X-ray, and microscopic images, were used to train a shared backbone on classification, segmentation, and detection tasks. Currently, the impact of tasks outside the histopathology domain remains unclear due to the diverse pre-training of the encoder. This impact on performance and robustness requires further investigation.
The following sections focus in more detail on two topics discussed in this paper. Although foundation models are still largely unexplored in terms of their application and performance, some approaches are mentioned below.
@@ 2.1 Foundation Models
[wang_transformer-based_2022] used data from the TCGA in combination with data from the pathology AI platform, PAIP, to train a modified swin transformer, called CTransPath (CTP), using self-supervision on 15 million patches. They presented an adapted contrastive loss, based on MoCo v3 [chen_empirical_2021], which uses a memory bank to retrieve the top S semantically relevant entries. These entries were used as additional positive examples for the loss calculation. The authors evaluated their model using patch classification, image retrieval, and weakly labeled WSI classification. Due to the large number of training images and the slow convergence of self-supervised training, they reported a training time of 250 hours on 48 GPUs (12.000 GPU-hours).
[campanella_computational_2023], presented a comparable approach, training a tiny vision transformer (ViT) using standard DINO, and a ViT base model using standard masked autoencoding (MAE), both trained on about 3 billion patches. The authors report a training time of over 3000 GPU-hours for the models that were evaluated on a variety of tasks ranging from disease detection to outcome prediction. The evaluation also included images scanned at a different hospital from the training slides.
[chen_towards_2024] present a general purpose foundation model that leverages over 100 million patches from 100.000 WSI slides across 20 major cancer types. They train a large ViT on patches, that were collected in an internal dataset, using DINOv2 [oquab2023dinov2]. They evaluate the model on 34 Tasks and find that it surpasses the previous baselines on most of them. The model was trained on 24 80GB GPUs.
Overall, all of the presented models rely on large image databases and require long training times, which contributes to increased CO2 emissions. The presented TC encoder and MTL training aim to reduce the need for large amounts of data while maintaining the desired performance. In addition, cross-center evaluation is needed to accurately asses models’ performances.
@@ 2.2 Weakly Labeled WSI Classification
Learning from WSIs that are only labeled on a case basis, or that have only one endpoint, is challenging because training on the entire image at once typically exceeds the GPU memory. In addition, since a WSI provides only one sample, many WSIs are needed to effectively train a deep learning model. Classification of such gigapixel images is therefore typically performed using multiple instance learning (MIL) [campanella_clinical-grade_2019, ghaffari_laleh_benchmarking_2022]. Using MIL involves two parts: first, extracting features from patches of the WSI using a pre-trained encoder to convert them into their latent representations, and second, aggregating features from a WSI using a trainable MIL head to predict the given label [liu_advmil_2024]. Therefore, robust encoders are needed to obtain patch representations that facilitate the second step of MIL [godson_immune_2024]. In this paper, MIL is used as an evaluation procedure to test the representativeness of the encoder’s features. The following presents the most commonly used approaches that focus on solving the second stage of MIL using either attention or convolution-based methods.
[lu_data-efficient_2021] introduced CLAM, a clustering-constrained attention MIL algorithm. The authors trained an attention-based head on features extracted from patches of a WSI to classify the corresponding labels. The attention was then used to identify subregions of high diagnostic value, which in turn were used to classify the entire slide. In addition, instance-level clustering was applied over the representative regions to constrain and refine the feature space.
@@ 3.2 Encoder Training
Since transformer-based approaches are currently dominant in the literature we chose a tiny swin transformer as shared encoder [liu_swin_2021, azad_advances_2024] (further denoted as TC-Swin). The encoder shares the same latent dimension with CTransPath, 768, and has 27.5M trainable parameters. Additionally, we compared the attention-based backbone to a tiny ConvNeXt architecture that shares a similar number of parameters and latent dimension but follows a convolution-based approach [liu_convnet_2022] (further denoted as TC-Conv). Both networks were initialized with the corresponding ImageNet-1k weights which were imported from the torchvision library [russakovsky_imagenet_2015]. During pre-training, the networks were optimized for 5 million steps until convergence, where one step was defined as computing the loss for one batch of one task. AdamW optimizer [loshchilov_decoupled_2019] was used with a constant learning rate of $10^{-4}$, a weight decay of 0.01, and a gradient accumulation of 128 steps. In total, the models were trained for 160 hours on one NVidia RTX A5000 GPU and the final models were used for evaluation.
@@ 4.2 Evaluation datasets
The learned representations of Tissue Concepts were evaluated using weakly labeled WSI classification tasks. Breast, prostate, and caolorectal tissue represent tissues within the training domain, lung cancer tissue was deliberately chosen to represent tissues outside the training domain. The architecture and learning schedule as described in Section 4.3, was used during the training of each MIL head. Where possible, multi-site datasets were split so that evaluation sites were not included in the training. The feature extractors were evaluated on the following datasets for each organ:
@@
Prostate Cancer: The Prostate cANcer graDe Assessment (PANDA) challenge provides 10616 openly available WSIs from two different centers, Karolinska and Radboud [bulten_artificial_2022]. Each of the WSIs was assigned an ISUP score resulting from the primary and secondary Gleason pattern. The score ranges from 0 to 5, with 0 representing normal tissue and higher numbers representing more severe prostate cancer. This information was used to create binary classification task or evaluation, where the task consists of differentiating between normal (ISUP 0) and cancerous tissue (ISUP $>$ 0). To test cross-center performance and overall performance, three different experiments were created from this dataset. First, the entire set of images obtained from the Radboud UMC was used as both the training and validation set, using a split ratio of 90% for training and 10% for validation. All images provided by the Karolinska Institute were used as the hold-out test set. In a subsequent experiment, the roles of training and test centers were swapped. In addition, a 5-fold cross-validation experiment involving both centers was conducted. We note that a small percentage of the dataset in the form of pre-selected patches were used during pre-training of the encoder, however, the task-specific head was discarded. To evaluate the effect of possible data leakage [kapoor_leakage_2023] an additional encoder was trained without the PANDA dataset in pre-training and evaluated only on this task. [mahdi_behzadi_weakly-supervised_2022] used this dataset in a k-fold cross-validation to evaluate the performance of their model and reported the accuracy and area under the curve for binary classification. Their reported results serve as an external baseline. [mahdi_behzadi_weakly-supervised_2022] reported the accuracy and AUC of their algorithm on the dataset. To be comparable, we also report these metrics in our experiments. Additionally, we report the F1-score as per-class metric, as suggested by [maier-hein_metrics_2024].
@@
Colorectal Cancer: During the SemiCOL challenge, a dataset of 499 WSIs from 4 different centers was provided as training dataset. The challenge task was to predict whether a given slide contained cancerous tissue or normal tissue. During the challenge, we used a modified model presented in [schafer_overcoming_2023] as an initial starting point and further fine-tuned on several colon and colorectal tasks. Using the learned encoder to extract features from the WSI, a MIL head was successively trained on these features. This resulted in an external test AUC of 0.99. In the experiments presented in this paper, the challenge training data was split by slide provider, effectively creating four smaller subsets of the original challenge dataset. For the evaluation presented in this paper, four different experiments were conducted. Images from one center were used as training examples (90/10), while images from the other three centers served as the hold out test set. By treating each center as training provider once, and using the other three centers as test centers, four different configurations were created. We note that a small percentage of patch data of the dataset was used during pre-training of the encoder, however, the task-specific head was discarded. To the best of our knowledge, there is no other publication reporting results on this specific dataset, so we are unable to report an external baseline. Since the dataset is balanced with regard to the class distribution for each center in all splits, we reported the Accuracy, F1-Score and AUC as metrics, following [maier-hein_metrics_2024].
@@ 4.3 Training of the MIL head
To aggregate the latent WSIs, a small CNN with global max pooling, was used. The initial layer utilized a 1x1 convolution, reducing the latent WSIs’ 768 channel to 16. Following this, 3x3 convolutions with a stride of 2, and padding of 1 doubling in depth each time were applied. Instance normalization, leaky ReLU activation, and 10% 2D dropout were applied in the convolutional layers. The lowest feature pyramid of 64 channels was globally max pooled, followed by one hidden linear layer and a classification layer. This resulted in 19250 total trainable parameters. The convolutional layers were initialized by drawing from a Kaiming normal distribution with a fixed seed. Additionally, we tested an attention-based approach by replacing the max pooling layer with an attention pooling layer, as suggested by [ilse2018attention]. Overall, we found similar results with both approaches across all organs. The results using an ABMIL aggregation head can be found in the subsection B.1.
@@ 5.2 Prostate Cancer
Table 3: Prostate: Tissue-specific model performance when applied to the PANDA challenge data. The model consist of organ-specific head and frozen encoders: ImageNet, CTP, TC-Swin, TC-Conv, TC-Swin-NoPanda. The table shows mean weighted Accuracy, F1 score, AUC, and standard deviations over 5 different runs for 5-fold cross-validation and cross-center transfer performance. The best score is shown in bold.
@@
TC-Swin-NoPanda 0.821 $\pm$ 0.07 0.807 $\pm$ 0.06 0.875 $\pm$ 0.007
Table 3 shows the results obtained of five runs on the PANDA dataset with cross-center and mixed evaluation. The ’clinic’ column indicates different settings, where "Karolinska $\rightarrow$ Radboud" refers to the MIL head being trained on data from the Karolinska Institute and evaluated on data from the Radboud Clinic. The additional encoder TC-Swin-NoPanda was trained without PANDA data in the pretraining.
When comparing the results obtained in the 5-fold cross-validation with the external baseline, both TC encoders show performances that were close to the external baseline, with TC-Conv slightly surpassing it with a mean accuracy of 0.957. The mean accuracies of CTP and ImageNet are 2 and 5 percentage points below the selected baseline, respectively. All foundation models outperform the pure ImageNet encoder by at least three percentage points. In terms of accuracy and AUC scores, two TC encoders slightly outperform the CTP weights by one percentage point with mean acc of 0.936 and mean auc of 0.979. When training the TC-Swin encoder without the PANDA dataset, a slight drop in performance can be observed, resulting in similar metrics compared to the CTP encoder.
In the cross-center evaluation, the performance of all models decreases compared to an in-center evaluation. CTP achieves a mean accuracy of 0.819 when transferring decision boundaries from the Karolinska site to Radboud subset. The opposite scenario results in a decrease of 14 percentage points (0.678 mean acc). Furthermore, an even steeper drop in mean accuracy and AUC from 5-fold cross-validation to cross-center validation is observed when using the ImageNet weights (from 0.900 to 0.658). All TC encoders are capable of extracting robust features. The TC-Swin encoder (from 0.942 to 0.846) shows more stable performance than the TC-Conv encoder (from 0.957 to 0.764) when compared to the metrics reported in the cross-center evaluation. The TC-Swin encoder trained without panda shows slightly lower performance on average to the TC-Swin encoder trained with PANDA. Independently of including the PANDA dataset in the pre-training, all TC-based encoders show a stronger cross-center performance over CTP.